OpenStudy (anonymous):

Hi can someone help me figure out what -10<-13+v is? I'm solving for the inequality of v. This is for a review! Thank you!

OpenStudy (texaschic101):

-10 < -13 + v ---- add 13 to both sides -10 + 13 < -13 + 13 + v ---- combine terms 3 < v any questions ?

OpenStudy (anonymous):

I had it right then! YAY I'm kind of just double checking my work right now :)

OpenStudy (texaschic101):

no problem with that...confirmation is good :)

OpenStudy (anonymous):

I'm a fan of you! I may have a few more questions I'd like to double check...hopefully that will be okay? I always get really nervous that people think I'm cheating on this site or something...

OpenStudy (texaschic101):

I will check the ones I know how to do...lol. I am glad to help when I can because sometimes I might need help.

OpenStudy (anonymous):

1/8-1/2 = 5/8 which is equal to 3 right?

OpenStudy (anonymous):

oh wait shoot wrong sign

OpenStudy (anonymous):

I just caught myself

OpenStudy (anonymous):

answer is 3/8?

OpenStudy (texaschic101):

1/8 - 1/2 = 1/8 - 4/8 = -3/8

OpenStudy (texaschic101):

understand ?

OpenStudy (anonymous):

is it really -3/8? or just 38 though? cause aren't you just subtracting I don't think it's a negative

OpenStudy (anonymous):

3/8*

OpenStudy (texaschic101):

1/8 - 1/2 ....common denominator is 8 1/8 - 4/8 = -3/8 (100% positive) The reason that it is negative is because 1/2 (or 4/8) is bigger then 1/8. Its like saying 3 - 5 = -2.....5 is bigger, and when you subtract a bigger number from a smaller number, it will be negative

OpenStudy (anonymous):

oh so if it was 1/2-1/8 THEN it would be positive 3/8, yes?

OpenStudy (texaschic101):

1/2 - 1/8 = 4/8 - 1/8 = 3/8.....yep

OpenStudy (anonymous):

I flat out can't do this one yet....-3/4 + 3/-5

OpenStudy (anonymous):

This probably seems really easy to you...people keep telling me it's really basic math...but I'm a writer and I truly suck at math -.-

OpenStudy (texaschic101):

-3/4 + (-3/5) .......3/-5 can be written as -3/5 -3/4 - 3/5 --- common denominator is 20 -15/20 - 12/20 = - 27/20 (when you have two negatives like this, you add and keep the same sign) answer is : -27/20, and if it needs to be in mixed number form, it is - 1 7/20

OpenStudy (anonymous):

It's supposed to be written in the simplest form which I think means keep it as a fraction and not as a decimal right?

OpenStudy (anonymous):

so this is written as -27/20 as in the negative sign is directly next to the 27 right?

OpenStudy (anonymous):

\[\frac{ -27}{ 20}\]

OpenStudy (texaschic101):

right..-27/20

OpenStudy (anonymous):

I think I'm good on the next three I have here...so how about this one?

OpenStudy (anonymous):

The Sugar Sweet Company is going to transport its sugar to market. It will cost to rent trucks, and it will cost an additional for each ton of sugar transported. Let C represent the total cost (in dollars), and let S represent the amount of sugar (in tons) transported. Write an equation relating C to S

OpenStudy (anonymous):

Oops, it will cost 4375 to rent the trucks and an additional 175 for each ton of sugar

OpenStudy (anonymous):

You know when I'm done here with you I'm gonna go to the English section to pay back all the help I got :) Not before I write you an awesome testimonial of course :D

OpenStudy (texaschic101):

C = 4375 + 175s

OpenStudy (anonymous):

right 4550

OpenStudy (texaschic101):

I wrote a little s only because a capital S can be confusing.....175S..see what I mean

OpenStudy (anonymous):

oh yeah! it looks like a 5

OpenStudy (texaschic101):

no..it is not 4550 it is : C = 4375 + 175s....you do not add the numbers

OpenStudy (anonymous):

But doesn't it say that C represents the total cost in dollars? wouldn't that be telling me to add them together?

OpenStudy (anonymous):

I have to plot this so can I actually use c and s in my equation?

OpenStudy (texaschic101):

total cost in dollars.....4375 is the cost to rent the trucks. And it costs 175 per ton of sugar. Your question is asking for an equation, not an exact answer. Because the answer can be different depending on the amount of sugar. Lets say you have 2 tons of sugar... C = 4375 + 175(2) C = 4375 + 350 C = 4725 but if you only had 1 ton of sugar.. C = 4375 + 175(1) C = 4375 + 175 C = 4550 The C will depend on how many tons of sugar. So the only thing you can do is set up the equation to solve for any value you sub in for s.

OpenStudy (anonymous):

ohhh that makes sense

OpenStudy (anonymous):

but does that mean I have to solve for s to get the equation I need to plot my points?

OpenStudy (anonymous):

you know you've helped me plenty! I can bring my questions up on the board for others to solve too :) I can see you're busy helping lots of other people. This doesn't seem fair